OpenStudy (anonymous):

A scale model of an office building is 3' x 2' x 5' (length, width, height). If the actual building has a length of 45 feet, what is the volume of the actual building? 81,000 cubic feet 102,150 cubic feet 101,250 cubic feet 30,000 cubic feet

OpenStudy (jadedry):

@ericmabry If 3' is the length of the model and 45' is the length of the building it means that the building is 45/3 = 15 times longer than the scale. Has an area (15^2) times larger than the scale. And a volume (15^3) times larger than the scale. You have two options: Convert all scale lengths to real lengths and multiply: \[(3*15) * (2*15) * (5*15) = 101 250 \] Or find the volume of the scale model: \[3*2*5 = 30\] \[30 * (15^3)\] =\[101 250\] either way you get the same answer.